# Harmonic covers of skeleta

### v1

@@ Abstract.
Primary 14G22, 31C20; secondary 14D10, 32P05, 14T20, 31C05
The author was supported by the Fund for Scientific Research Flanders (PhD Fellowship 11F0123N and Research Project G0B1721N)
@@ 1. Introduction
  • Thirdly we show that that finite covers of skeleta are harmonic morphisms (Theorem 1.22).
@@
Let us now give a more detailed overview of this work.
@@
Throughout we work over a discretely valued ground field $k$ whose ring of integers will be denoted by $k^{\circ}$. For simplicity the residue field $\tilde{k}$ is assumed algebraically closed. Log-structures are defined with respect to the Zariski site, and we equip $S=\mathrm{Spec}\,k^{\circ}$ with the “standard” log-structure $S^{\dagger}$ defined via $k^{\circ}\setminus 0\to k^{\circ}$.
In this article, a toroidal scheme over $S$ is defined as an integral separated finite type $S$-scheme ${X}$ equipped with a Zariski log-structure
@@
  • toric pairs: ${X}$ is a toric $S$-scheme, $D_{{X}}$ is torus-invariant (containing ${X}_{s}$).
@@
A toroidal $S$-scheme ${X}^{\dagger}$ is stratified by the intersections of components of $D_{{X}}$, also called the log-stratification. The dual complex $\mathcal{D}({X}^{\dagger})$ is obtained as the dual intersection complex of the log-stratification and is a polyhedral complex which may possibly have some unbounded polyhedra corresponding to horizontal log-strata. An important fact is that the dual complex $\mathcal{D}({X}^{\dagger})$ admits a natural topological embedding
@@
Throughout the paper, we denote by
@@
a finite morphism of proper toroidal $S^{\dagger}$-schemes or short a toroidal cover, and the associated cover of dual complexes is denoted by
@@
We are primarily interested in such covers because they provide an adequate “monomialised” setting to study analytifications of covers of $k$-varieties via degeneration techniques. There is an advantage of working with toroidal schemes as opposed to snc pairs: finite covers of snc pairs hardly arise in nature – the so-called “impossibility of simultaneous resolution” according to Abhyankar [Kat94, §12]. In contrast, if in the above only $f_{\eta}$ and ${X}^{\dagger}$ are given then the normalisation $f:X^{\prime\dagger}\to{X}^{\dagger}$ is a toroidal cover assuming $f$ is tamely ramified along $D_{X}$, and Kato raised the question whether in wildly ramified situations toroidal covers also exist (see §3.5).
@@
To simplify terminology we define a $\mathbb{Z}$-PL complex as a strictly convex rational polyhedral complex equipped with a $(\mathbb{Z},|k^{\times}|)$-piecewise linear structure (also known as an integral $k$-affine structure). In particular each polyhedron $\sigma$ of a $\mathbb{Z}$-PL complex $\Sigma$ is equipped, in a compatible fashion, with a lattice $M_{\sigma}$ of $\mathbb{Z}$-affine functions. For toroidal $S$-schemes ${X}^{\dagger}$ the dual complex $\mathcal{D}({X}^{\dagger})$ comes canonically equipped with a $\mathbb{Z}$-PL structure where the lattices $M_{\sigma}$ consist of toric coordinates on $\sigma$. More precisely every polyhedron $\sigma$ of $\mathcal{D}({X}^{\dagger})$ corresponds to a unique generic point $x$ of a log-stratum of ${X}^{\dagger}$ and the lattice of $\mathbb{Z}$-affine functions $M_{\sigma}$ is then given by
@@
(1.7) $M_{\Sigma}\coloneqq\lim_{\sigma\in\Sigma}M_{\sigma}$
1.8Local degrees and the lattice index.
@@
be a finite cover of $n$-dimensional $\mathbb{Z}$-PL complexes. Consider maps of a facet-ridge pair $(\tau^{\prime},\sigma^{\prime})$ of $\Sigma^{\prime}$ to a given facet-ridge pair $(\tau,\sigma)$ of $\Sigma$ – see Figure 1.11 for a picture. Then, fixing $\tau^{\prime}$, we call $\phi:\Sigma^{\prime}\to\Sigma$ balanced at $\tau^{\prime}$ if the following sum of lattice indices, called the local degree at $\tau^{\prime}$, is is independent of the choice of $\sigma$ and thus well-defined :
@@
Theorem 1.12 (Balancing formula, See Theorem 3.24).
@@
1.13Metrisation of $\mathbb{Z}$-PL complexes.
Any $\mathbb{Z}$-PL complex $\Sigma$ can be equipped with a canonical measure $\mu_{\Sigma}$ that we call the conformal measure, such that isometries are isomorphisms (i.e. unimodular bijections). In essence, it is the unique piecewise Lebesgue measure which satisfies the following conformality property (Lemma 4.7): for a finite cover of bounded $\mathbb{Z}$-PL polyhedra $\sigma^{\prime}\to\sigma$, we have
@@
Given a $\mathbb{Q}$-Cartier divisor $\mathscr{D}$ on ${X}$ we have intersection numbers $\mathscr{D}\cdot C\in\mathbb{Q}$ for any proper regular curve $C\subset{X}$. We define the specialisation $\mathrm{sp}_{*}\mathscr{D}$ as the combinatorial divisor
(1.16) $\mathrm{sp}_{*}\mathscr{D}\coloneqq\sum_{\tau}\frac{\mathscr{D}\cdot E_{\tau}}{\mathrm{vol}(\tau)}[\tau],$
summing over all vertical ridges $\tau$ of $\Sigma$; here $E_{\tau}$ is the $1$-dimensional log-stratum associated to $\tau$. In fact $\mathrm{sp}_{*}\mathscr{D}$ only depends on the linear equivalence class of $\mathscr{D}$.
1.17.
From the remainder of the introduction we assume that the dual complex $\mathcal{D}({X}^{\dagger})$ is simplicial, equivalently Weil divisors supported on $D_{X}$ are $\mathbb{Q}$-Cartier (5.8).
1.18Laplacians of PL functions.
For the definition of slopes of PL functions, we need to take into account additional structure next to the $\mathbb{Z}$-PL structure of the dual complex $\Sigma=\mathcal{D}({X}^{\dagger})$. In the spirit of the works [GRW16, Car20] we additionally decorate $\Sigma$ with some intersection numbers called $\alpha$-constants – this gives rise to the notion of a $\mathbb{Z}$-PL tropical complex (Definition 6.12). Informally speaking the slope $\partial_{\sigma/\tau}F$ of a $\mathbb{Z}$-affine PL function $F:\Sigma\to\mathbb{R}$ on a $\mathbb{Z}$-PL tropical complex $(\Sigma,\alpha)$ along a facet-ridge pair $(\sigma,\tau)$ should be thought of as an appropriate directional derivative of $F$ in a direction normal to $\tau$ that is distinguished by the $\alpha$-constants. The precise definition is given in 6.15.
@@
summing over facet-ridge pairs $(\sigma,\tau)$ with $\tau$ bounded; and we call $F$ harmonic if $\Delta(F)=0$.
@@
Similarly for $\mathbb{Q}$-Cartier divisors.
1.21.
By the slope formula (1.20) it follows that $F_{\mathscr{D}}$ is harmonic if and only if $\mathscr{D}$ is numerically trivial on the vertical $1$-dimensional log-strata. This is for instance the case if $\mathscr{D}$ is a principal divisor, and then the function $F_{\mathscr{D}}$ is usually called smooth.
@@
1.24An extension to compactified skeleta.
@@
1.25Reminders on Riemann-Hurwitz formula.
@@
$\chi(X)-(\deg g)\chi(Y)=\sum_{x\in X}\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}.$
The summand of the right-hand side is also known as the different $\delta_{g}(x)\coloneqq\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}$ of $g$ at $x$.
@@
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite seperable cover of proper toroidal $S$-schemes. Let $\phi:\Sigma^{\prime}\to\Sigma$ be the assocated cover of skeleta. Then there exists a $\mathbb{Z}$-PL function $\delta:\Sigma^{\prime}\to\mathbb{R}_{\geq 0}$, called the different function, such that for each divisorial point $x^{\prime}\in\Sigma^{\prime}$ and $x=\phi(x^{\prime})$ we have
@@
  • Suppose $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})$ is a nonzero holomorphic canonical form and let $\mathrm{wt}_{\omega}:{X}^{\mathrm{mon}}\to\mathbb{R}_{\geq 0}$ denote the associated weight function of [MN15]. Then
    $\Delta(\mathrm{wt}_{\omega}|_{\Sigma})=K_{\Sigma}-\rho_{*}\mathrm{div}\,\omega;$
@@
  • There exists a function $\lambda:X^{\mathrm{an}}\to\mathbb{R}_{\geq 0}$ which is $\mathbb{Z}$-affine on the faces of dual complexes, such that if $x\in X^{\mathrm{an}}$ is divisorial we have $\lambda(x)=\frac{1}{e(\mathscr{H}(x)/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/k^{\circ}}$; this gives some evidence for an expectation of Temkin on the piecewise linearity of Kähler norms (Remark 8.11).
1.29.
(Companion paper) The companion paper [Wae25] contains some complements and applications in the case of curves. There we study the metrisation of Berkovich curves over discretely valued fields in more detail and show that the different function controls simultaneous curve skeleta (shown earlier in [CTT16] with different hypotheses) and give applications to reduction types of arithmetic curves and wild surface quotient singularities.
1.30Related work.
The balancing formula of 1.12 is strongly analogous to tropical multiplicity formulas [ST08, BPR16, GRW16, GJR24], where similar sums of lattice indices play the role of the degree. Our result is different since we consider covers of a skeleton by a skeleton and not covers of a tropicalisation by a skeleton. Roughly speaking our case corresponds to the setting of tropical multiplicity one. A similar notion of a balanced cover of complexes in all dimensions appears in [Var24].
The notion of a harmonic morphism presented in 1.22 closely mimics the eponymous notion in Riemannian geometry [Fug78, Ish79, BW03, EF01], but is not directly related. In dimension one discrete analogues were well-studied, in particular see the works [Ura00, BN09, ABBR15].
@@
The main inspiration for the slope formula 1.19 is the prior work [GRW16], which deals with the strictly semistable case over algebraically closed ground fields and builds on earlier work in [Car20, BPR16].
The different function for covers of analytic curves was studied systematically in a series of works [Tem16, Tem17, CTT16, BT20] over algebraically closed ground fields, and has its historical origin in Gabber’s work on the $p$-adic Riemann existence problem (80s unpublished), see [Lüt93]. The Riemann-Hurwitz formula 1.26 for possibly wildly ramified covers of skeleta was pioneered in [CTT16] over curves over algebraically ground fields, see the references there for earlier work on wild ramification of analytic curves. In upcoming work, Hübner-Temkin [HT25] study analytic properties of the different function over any nonarchimedean field, but not the Riemann-Hurwitz formula.
After a preliminary version of this paper was written, Yanbo Fang kindly pointed the author to the earlier work of A. Gross [Gro18], where tropical intersection theory is developed on the dual complex of a toroidal variety over a field. There is some overlap, but his results do not adress mixed characteristic situations.
1.31Overview of the paper.
Section 2 covers preliminaries on logarithmic regularity. For convenience we give a self-contained exposition of the tools from logarithmic geometry that we use. We have put all necessary background on polyhedral geometry and fans in a separate Appendix A. In Section 3 we review dual complexes of toroidal $S$-schemes, describe their $\mathbb{Z}$-PL structure and prove Theorem 1.12. In Section 4 we construct the conformal measures on $\mathbb{Z}$-PL complexes and give explicit descriptions for dual complexes. Section 5 studies specialisation of Cartier divisors and some intersection numbers are computed. Section 6 introduces slopes and the Laplacian of PL functions and contains the proofs of Theorems 1.19 and  1.22. Section 7 contains some preparations on logarithmic differentials, our comparison of canonical sheaves and log-canonical sheaves may be of independent interest. In Section 8 we discuss potential theory of differential forms and prove Theorem 1.26.
1.32Conventions and notation.
@@
Throughout we denote by $k$ a discretely valued field. Its ring of integers is denoted by $k^{\circ}$ and its residue field, assumed algebraically closed, is denoted by $\tilde{k}$. We let $p\geq 1$ denote the characteristic exponent of $\tilde{k}$. We write $\varpi$ for a choice of uniformiser of $k^{\circ}$. We write $S=\mathrm{Spec}\,k^{\circ}$. The generic and closed point of $S$ are denoted by $\eta$ and $s$ respectively, so $S=\{\eta,s\}$ as a set.
@@
For the reader’s convenience, we discuss conventions on monoids, cones, fans, polyhedra, and $\mathbb{Z}$-PL spaces at length in appendix A, in summary:
@@
  • Given a polyhedral complex $\Sigma$ we write $\Sigma(n)$ for the set of $n$-dimensional polyhedrons of $\Sigma$.
@@
  • If $\mathscr{M}$ is a monoid, we write $\mathscr{M}^{\times}$ for its group of invertible elements, $\mathscr{M}^{\sharp}\coloneqq\mathscr{M}/\mathscr{M}^{\times}$ for its sharpification and $\mathscr{M}^{\mathrm{gp}}$ for its groupification. We denote by $\mathscr{M}^{\mathrm{int}}$ and $\mathscr{M}^{\mathrm{sat}}$ the integralisation and saturation respectively. Most monoids we encounter will be fs, i.e. fine and saturated.
Berkovich $k$-analytification $(\cdot)^{\mathrm{an}}$ will only be used in Section 8. Completed residue fields are denoted $\mathscr{H}(\cdot)$.
@@
1.33Acknowledgements.
The results of this article are part of the author’s PhD thesis, supervised by Johannes Nicaise; I am very grateful for his guidance and advice over the years. It is also my pleasure to thank Stefan Wewers, Lars Halvard Halle, Michaël Maex, Michael Tëmkin, Yanbo Fang, Antoine Ducros, Alejandro Vargas and Soham Karwa for valuable discussions related to this work.

### v2

@@ Abstract.
Primary 14G22, 31C20; secondary 14D10, 14T20, 31C05
The author was supported by the Fund for Scientific Research (FWO) Flanders, grants 11F0123N and G0B1721N. The results in this paper are part of my PhD thesis supervised by Johannes Nicaise, I am very grateful for his guidance and advice over the years. I also thank Stefan Wewers, Lars Halvard Halle, Michael Temkin, Yanbo Fang, Antoine Ducros, Michaël Maex, Alejandro Vargas and Soham Karwa for valuable discussions.
@@ 1. Introduction
  • Thirdly we show that finite covers of skeleta are harmonic morphisms (Theorem 1.21).
@@
Let us now give an overview of this work.
@@
Throughout we work over a discretely valued ground field $k$ whose ring of integers will be denoted by $k^{\circ}$. The residue field $\tilde{k}$ is assumed algebraically closed. Log-structures are defined with respect to the Zariski site, and we equip $S=\mathrm{Spec}\,k^{\circ}$ with the “standard” log-structure $S^{\dagger}$ defined via $k^{\circ}\setminus 0\to k^{\circ}$.
In this paper a toroidal scheme over $S$ is defined as an integral separated finite type $S$-scheme ${X}$ equipped with a Zariski log-structure
@@
  • toric pairs: ${X}$ is a toric $S$-scheme andµ $D_{{X}}$ is torus-invariant (containing ${X}_{s}$).
@@
A toroidal $S$-scheme ${X}^{\dagger}$ is stratified by the intersections of components of $D_{{X}}$, also called the log-stratification. The dual complex $\mathcal{D}({X}^{\dagger})$ is obtained as the dual intersection complex of the log-stratification and is a polyhedral complex which may possibly have some unbounded polyhedra corresponding to horizontal log-strata. An important fact is that the dual complex $\mathcal{D}({X}^{\dagger})$ admits a canonical topological embedding
@@
Throughout the paper we denote by
@@
a finite morphism of toroidal $S^{\dagger}$-schemes or short a toroidal cover, and the associated cover of dual complexes is denoted by
@@
We are primarily interested in such covers because they provide an adequate “monomialised” setting to study analytifications of covers of $k$-varieties via degeneration techniques. There is an advantage of working with toroidal schemes as opposed to snc pairs: finite covers of snc pairs hardly arise in nature – the so-called “impossibility of simultaneous resolution” according to Abhyankar [Kat94, §12]; also see the negative results of [LL99]. In contrast, if in the above only $f_{\eta}$ and ${X}^{\dagger}$ are given then the normalisation $f:X^{\prime\dagger}\to{X}^{\dagger}$ is a toroidal cover assuming $f$ is tamely ramified along $D_{X}$. Kato raised the question whether in wildly ramified situations toroidal covers also exist (see §3.5).
@@
In this paper $\mathbb{Z}$-PL complex is defined as an integral $k$-affine strictly convex rational polyhedral complex. Roughly speaking a $\mathbb{Z}$-PL complex is glued from polyhedra defined by systems of inequalities of the form $a_{1}x_{1}+\dots+a_{n}x_{n}\geq b$ where $b,a_{1},\dots,a_{n}\in\mathbb{Z}$. Each polyhedron $\sigma$ of a $\mathbb{Z}$-PL complex $\Sigma$ is equipped with a lattice $M_{\sigma}$ of $\mathbb{Z}$-affine functions and for each pair of faces $\tau\leq\sigma$ we have compatible restriction maps $M_{\sigma}\to M_{\tau}$. For toroidal $S$-schemes ${X}^{\dagger}$ the dual complex $\mathcal{D}({X}^{\dagger})$ comes canonically equipped with a $\mathbb{Z}$-PL structure where the lattices $M_{\sigma}$ consist of toric coordinates on $\sigma$. More precisely every polyhedron $\sigma$ of $\mathcal{D}({X}^{\dagger})$ corresponds to a unique generic point $x$ of a log-stratum of ${X}^{\dagger}$ and the lattice of $\mathbb{Z}$-affine functions $M_{\sigma}$ is then given by
@@
$M_{\Sigma}\coloneqq\lim_{\sigma\in\Sigma}M_{\sigma}.$
1.7Local degrees and the lattice index.
@@
be a finite cover of $n$-dimensional $\mathbb{Z}$-PL complexes. Consider maps of a facet-ridge pair $(\tau^{\prime},\sigma^{\prime})$ of $\Sigma^{\prime}$ to a given facet-ridge pair $(\tau,\sigma)$ of $\Sigma$ – see Figure 1.10 for a picture. Then, fixing $\tau^{\prime}$, we call $\phi:\Sigma^{\prime}\to\Sigma$ balanced at $\tau^{\prime}$ if the following sum of lattice indices, called the local degree at $\tau^{\prime}$, is independent of the choice of $\sigma$ and thus well-defined :
@@
Theorem 1.11 (Balancing condition, See Theorem 3.24).
@@
1.12Metrisation of $\mathbb{Z}$-PL complexes.
Any $\mathbb{Z}$-PL complex $\Sigma$ can be equipped with a canonical piecewise Lebesgue measure $\mu_{\Sigma}$ that we call the conformal measure such that isometries equal isomorphisms (i.e. unimodular bijections). It is determined by the following conformality property, see Lemma 4.7: for any finite cover of bounded $\mathbb{Z}$-PL polyhedra $\sigma^{\prime}\to\sigma$, we have
@@
1.14.
For the remainder of the introduction we consider proper toroidal $S$-schemes ${X}$ for which the dual complex $\mathcal{D}({X}^{\dagger})$ is simplicial, equivalently Weil divisors supported on $D_{X}$ are $\mathbb{Q}$-Cartier (5.8).
@@
Given a $\mathbb{Q}$-Cartier divisor $\mathscr{D}$ on ${X}$ which is supported on $D_{X}$ we have intersection numbers $\mathscr{D}\cdot C\in\mathbb{Q}$ for any proper regular curve $C\subset{X}$. We define the specialisation $\mathrm{sp}_{*}\mathscr{D}$ as the combinatorial divisor
(1.16) $\mathrm{sp}_{*}\mathscr{D}\coloneqq\sum_{\tau}\frac{\mathscr{D}\cdot E_{\tau}}{\mathrm{vol}(\tau^{-})}[\tau],$
summing over all vertical ridges $\tau$ of $\Sigma$; here $E_{\tau}$ is the $1$-dimensional log-stratum associated to $\tau$ and $\tau^{-}$ is a bounded subset of $\tau$ (cf. 4.21). In fact $\mathrm{sp}_{*}\mathscr{D}$ only depends on the linear equivalence class of $\mathscr{D}$.
1.17Laplacians of PL functions.
For the definition of slopes of PL functions, we need to take into account additional structure next to the $\mathbb{Z}$-PL structure of the dual complex $\Sigma=\mathcal{D}({X}^{\dagger})$. In the spirit of the works [GRW16, Car20] we additionally decorate $\Sigma$ with some intersection numbers called $\alpha$-constants – this gives rise to the notion of a $\mathbb{Z}$-PL tropical complex (Definition 6.12). Informally speaking the slope $\partial_{\sigma/\tau}F$ of a $\mathbb{Z}$-affine PL function $F:\Sigma\to\mathbb{R}$ on a $\mathbb{Z}$-PL tropical complex $(\Sigma,\alpha)$ along a facet-ridge pair $(\sigma,\tau)$ should be thought of as an appropriate directional derivative of $F$ in a direction normal to $\tau$ that is distinguished by the $\alpha$-constants. . The precise definition is given in 6.15.
@@
summing over facet-ridge pairs $(\sigma,\tau)$ with $\tau$ vertical; and we call $F$ harmonic if $\Delta(F)=0$.
@@
1.20.
By the slope formula (1.19) it follows that $F_{\mathscr{D}}$ is harmonic if and only if $\mathscr{D}$ is numerically trivial on the vertical $1$-dimensional log-strata. This is for instance the case if $\mathscr{D}$ is a principal divisor  [^1] In this case the associated functions $F_{\mathscr{D}}$ are usually called smooth in the literature..
@@
1.23An extension to compactified skeleta.
@@
1.24Reminders on Riemann-Hurwitz formula.
@@
$-\chi(X)+(\deg g)\chi(Y)=\deg\omega_{X/Y}=\sum_{x\in X}\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}.$
We also write $\delta_{g}(x)\coloneqq\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}$ for the different of $g$ at $x$.
@@
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite seperable cover of proper toroidal $S$-schemes with simplicial dual complexes. Denote $\phi:\Sigma^{\prime}\to\Sigma$ for the assocated cover of skeleta. Then there exists a $\mathbb{Z}$-affine function $\delta:\Sigma^{\prime}\to\mathbb{R}_{\geq 0}$, called the different function, such that for each divisorial point $x^{\prime}\in\Sigma^{\prime}$ and $x=\phi(x^{\prime})$ we have
@@
  • Suppose $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})$ is a nonzero holomorphic canonical form and let $\mathrm{wt}_{\omega}:\Sigma\to\mathbb{R}_{\geq 0}$ denote the associated weight function of [MN15]. Then
    $\Delta(\mathrm{wt}_{\omega})=K_{\Sigma}-\rho_{*}\mathrm{div}\,\omega;$
@@
  • The function $\lambda:X^{\mathrm{div}\,}\to\mathbb{R}_{\geq 0}x\mapsto\frac{1}{e(\mathscr{H}(x)/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/k^{\circ}}$ extends to a $\mathbb{Z}$-affine function on the faces of dual complexes. This gives some evidence for an expectation of Temkin on the piecewise linearity of Kähler norms (Remark 8.11).
1.28.
(Companion paper) The companion paper [Wae25] contains some examples in the case of curves and gives applications to reduction types of arithmetic curves.
1.29Related work.
The balancing formula of 1.11 is strongly analogous to tropical multiplicity formulas [ST08, BPR16, GRW16, GJR24], where similar sums of lattice indices play the role of the degree. Our result is different since we consider covers of a skeleton by a skeleton and not covers of a tropicalisation by a skeleton. Roughly speaking our case corresponds to the setting of tropical multiplicity one. A similar notion of a balanced cover of complexes in all dimensions appears in [Var24].
The notion of a harmonic morphism presented in 1.21 closely mimics the eponymous notion in Riemannian geometry [Fug78, Ish79, BW03, EF01], but is not directly related. In dimension one discrete analogues were well-studied, see the works [Ura00, BN09, ABBR15].
@@
The main inspiration for the slope formula 1.18 is the prior work [GRW16], which deals with the strictly semistable case over algebraically closed ground fields and builds on earlier work in [Car20, BPR16]. After the paper was written, Yanbo Fang kindly showed the work [Gro18], where tropical intersection theory is developed on the dual complex of a toroidal variety over a field. There is some overlap, but his results do not adress mixed characteristic situations. It is an interesting question whether his results on tropical chow groups carry over. Similarly one can wonder how the Poincaré-Lelong formula presented here relates to the ones developed in [Thu05, CD12, GJR24, GK17, Mih24], and to the theory of tropical intersection theory.
The different function and skeletal Riemann-Hurwitz formula was introduced and studied systematically in a series of works [Tem16, Tem17, CTT16, BT20] for covers of curves over algebraically closed ground fields, see the references cited in these papers for earlier work on wild ramification of analytic curves. In parallel upcoming work Hübner-Temkin [HT25] study analytic properties of the different function over any nonarchimedean field, but not the Riemann-Hurwitz formula.
1.30Overview of the paper.
Section 2 covers preliminaries on logarithmic regularity. For convenience we give a self-contained exposition of the tools from logarithmic geometry that we use. We have put all necessary background on polyhedral geometry and fans in a separate Appendix A. In Section 3 we review dual complexes of toroidal $S$-schemes, describe their $\mathbb{Z}$-PL structure and prove Theorem 1.11. In Section 4.8 we construct the conformal measures on $\mathbb{Z}$-PL complexes and give explicit descriptions for dual complexes. Section 5 studies specialisation of Cartier divisors and some intersection numbers are computed. Section 6 introduces slopes and the Laplacian of PL functions and contains the proofs of Theorems 1.18 and  1.21. Section 7 contains some preparations on logarithmic differentials, our comparison of canonical sheaves and log-canonical sheaves may be of independent interest. In Section 8 we discuss potential theory of differential forms and prove Theorem 1.25.
1.31Conventions and notation.
@@
Throughout we denote by $k$ a discretely valued field. Its ring of integers is denoted by $k^{\circ}$ and its residue field, assumed algebraically closed, is denoted by $\tilde{k}$. We let $p\geq 1$ denote the characteristic exponent of $\tilde{k}$ and $\varpi$ a choice of uniformiser of $k^{\circ}$. We write $S=\mathrm{Spec}\,k^{\circ}$. The generic and closed point of $S$ are denoted by $\eta$ and $s$ respectively, so $S=\{\eta,s\}$ as a set.
@@
Conventions on monoids, cones, fans, polyhedra, and $\mathbb{Z}$-PL spaces are discussed in appendix A, in summary:
@@
  • Given a polyhedral complex $\Sigma$ we write $\Sigma(n)$ for the set of $n$-dimensional polyhedra of $\Sigma$.
@@
  • If $\mathscr{M}$ is a monoid, we write $\mathscr{M}^{\times}$ for its group of invertible elements, $\mathscr{M}^{\sharp}\coloneqq\mathscr{M}/\mathscr{M}^{\times}$ for its sharpification and $\mathscr{M}^{\mathrm{gp}}$ for its groupification. Most monoids we encounter will be fs, i.e. fine and saturated.
Berkovich $k$-analytification is denoted by $(\cdot)^{\mathrm{an}}$; it will only be used in the end of Section 5 and in Section 8. Completed residue fields are denoted by $\mathscr{H}(\cdot)$.